OpenStudy (anonymous):

A dvd player was priced at 90$ and it was sold at a discount of 20%.After the discount,the shopkeeper still make a profit of 20%.what is the cost price of the dvd player? pls check if is my working and answer is wether correct or not.

OpenStudy (anonymous):

=100-20/100 times 90$ =72$ cost price+20%=72$ =72$/1 divide 120/100 =60$

OpenStudy (anonymous):

@MrNood

OpenStudy (mrnood):

Yes - that is correct

OpenStudy (anonymous):

ok.thanks @MrNood